# 22affiliationtext: Booth School of Business

### v1

@@ D.1.3 Regularization (MNIST)
A description of the architecture and experiment is given in Appendix C. Tables 7 and 8 respectively show results of group lasso (Yuan and Lin, 2006) and $\ell_{2}$ regularization, while Table 9 shows the effects regularizing the latent space of the VAE by varying $\beta$ (Higgins et al., 2017). [^2] All other experiments in Appendix D.1 are run with $\lambda=0$ for group lasso and $\ell_{2}$ and with $\beta=1$. Altogether, these results indicate relatively stable performance across regularization weights while suggesting that tuning these regularizers could be worthwhile (e.g., notice that $\lambda=10$ improves performance).
@@
Table 7: Group lasso regularization on MNIST
Metric / Evaluation $\lambda$=1 $\lambda$=10 $\lambda$=100
Validation Metrics
Validation ELBO (bpd) ↓ 0.2983 (0.0053) 0.2907 (0.0081) 0.2869 (0.0064)
Reconstruction Loss (bpd) ↓ 0.2683 (0.0047) 0.2605 (0.0088) 0.2568 (0.0074)
Reconstructed Distribution Evaluation
Covariance Error ↓ 1.2976 (0.0608) 1.2175 (0.0611) 1.3087 (0.0469)
Mean Error ↓ 0.0196 (0.0020) 0.0134 (0.0018) 0.0178 (0.0023)
MMD ↓ 0.0125 (0.0018) 0.0069 (0.0023) 0.0109 (0.0015)
Covariance Error (ivn) ↓ 2.2654 (0.2437) 2.2651 (0.2638) 2.2818 (0.2705)
Mean Error (ivn) ↓ 0.0277 (0.0018) 0.0277 (0.0025) 0.0315 (0.0031)
MMD (ivn) ↓ 0.0993 (0.0367) 0.0982 (0.0356) 0.0986 (0.0358)
Generated Distribution Evaluation
MMD ↓ 0.0048 (0.0005) 0.0053 (0.0005) 0.0057 (0.0009)
MMD (ivn) ↓ 0.0283 (0.0101) 0.0274 (0.0076) 0.0211 (0.0061)
Table 8: $\ell_{2}$ regularization on MNIST
Metric / Evaluation $\lambda$=1 $\lambda$=10 $\lambda$=100
Validation Metrics
Validation ELBO (bpd) ↓ 0.3079 (0.0138) 0.2917 (0.0063) 0.2966 (0.0084)
Reconstruction Loss (bpd) ↓ 0.2774 (0.0145) 0.2612 (0.0067) 0.2667 (0.0092)
Reconstructed Distribution Evaluation
Covariance Error ↓ 1.3562 (0.1205) 1.2883 (0.0103) 1.4347 (0.0955)
Mean Error ↓ 0.0178 (0.0021) 0.0182 (0.0018) 0.0209 (0.0023)
MMD ↓ 0.0090 (0.0020) 0.0088 (0.0022) 0.0119 (0.0025)
Covariance Error (ivn) ↓ 2.3289 (0.2539) 2.2137 (0.2341) 2.2341 (0.2606)
Mean Error (ivn) ↓ 0.0287 (0.0022) 0.0281 (0.0021) 0.0298 (0.0031)
MMD (ivn) ↓ 0.1023 (0.0377) 0.0957 (0.0348) 0.0881 (0.0317)
Generated Distribution Evaluation
MMD ↓ 0.0051 (0.0005) 0.0056 (0.0006) 0.0067 (0.0012)
MMD (ivn) ↓ 0.0231 (0.0071) 0.0254 (0.0072) 0.0217 (0.0064)

### v2

@@ E.1.3 Regularization (MNIST)
A description of the architecture and experiment is given in Appendix D. Tables 6 and 7 respectively show results of group lasso (Yuan and Lin, 2006) and $L_{2}$ regularization, while Table 8 shows the effects regularizing the latent space of the VAE by varying $\beta$ (Higgins et al., 2017). [^1] All other MNIST experiments in Appendix E.1 are run with $\lambda=0$ for group lasso and $L_{2}$ and with $\beta=1$. Altogether, the results are relatively stable performance across regularization strengths, with slight increase in reconstruction performance for higher group lasso or $L_{2}$ regularization strength but no other statistically significant differences.
@@
Table 6: Evaluation of CM (end-to-end) in terms of reconstruction and concept learning on MNIST for different group lasso regularization strengths.
$\lambda=$ 0 $\lambda=$1 $\lambda=$10 $\lambda=$100
reconstruction (ELBO BPD)
in-distribution 0.259 ± 0.005 0.259 ± 0.003 0.253 ± 0.003 0.248 ± 0.002
out-of-distribution
concept learning (sliced Wasserstein)
obs 0.099 ± 0.006 0.092 ± 0.005 0.091 ± 0.006 0.092 ± 0.006
scaled 0.062 ± 0.002 0.069 ± 0.002 0.075 ± 0.003 0.118 ± 0.006
shear 0.119 ± 0.006 0.118 ± 0.005 0.115 ± 0.008 0.113 ± 0.004
shift 0.113 ± 0.005 0.112 ± 0.007 0.116 ± 0.005 0.113 ± 0.003
swel 0.126 ± 0.007 0.122 ± 0.004 0.131 ± 0.007 0.127 ± 0.006
thic 0.178 ± 0.007 0.183 ± 0.008 0.216 ± 0.015 0.220 ± 0.005
thin 0.064 ± 0.004 0.072 ± 0.005 0.082 ± 0.009 0.111 ± 0.008
concept learning mean 0.109 ± 0.006 0.110 ± 0.005 0.118 ± 0.008 0.128 ± 0.005
Table 7: Evaluation of CM (end-to-end) in terms of reconstruction and concept learning on MNIST for different $L_{2}$-norm regularization strengths.
$\lambda=$ 0 $\lambda=$1 $\lambda=$10 $\lambda=$100
reconstruction (ELBO BPD)
in-distribution 0.259 ± 0.005 0.254 ± 0.003 0.255 ± 0.004 0.253 ± 0.001
out-of-distribution
concept learning (sliced Wasserstein)
obs 0.099 ± 0.006 0.091 ± 0.006 0.102 ± 0.005 0.086 ± 0.004
scaled 0.062 ± 0.002 0.068 ± 0.005 0.061 ± 0.003 0.073 ± 0.004
shear 0.119 ± 0.006 0.119 ± 0.007 0.127 ± 0.005 0.111 ± 0.006
shift 0.113 ± 0.005 0.111 ± 0.006 0.121 ± 0.005 0.108 ± 0.003
swel 0.126 ± 0.007 0.121 ± 0.009 0.133 ± 0.005 0.115 ± 0.004
thic 0.178 ± 0.007 0.186 ± 0.015 0.202 ± 0.010 0.183 ± 0.005
thin 0.064 ± 0.004 0.072 ± 0.008 0.062 ± 0.003 0.074 ± 0.006
concept learning mean 0.109 ± 0.006 0.110 ± 0.008 0.115 ± 0.005 0.107 ± 0.005